# Measurement-Induced Multimode Squeezed Light Interferometers with Scalable Architectures

### v1

@@ II Concept and theory
The expressibility of the unitary depends on several factors: the number of modes, the architecture of the graph (i.e. the correlation complexity) and the measurement strategy. The number of modes in the state can be easily controlled in the time domain, while the complexity of the graph state correlations can be varied by increasing the number of cascaded interferometers in the setup as shown in Fig. 1. Regarding the measurement strategy, we have so far assumed that spatial mode A was measured (as illustrated in Fig. 2a), thereby transforming the nodes in mode B. However, alternative measurement schemes can be employed, considering that with a fast optical switch each spatio-temporal mode could be directed either to the variable quadrature measurement or to the output. For example, starting from a one-dimensional graph (using a single interferometer, $K_{1}=1$), we find that the expressibility is slightly enhanced by employing a “Knight’s distance” measurement strategy, where the output modes alternate between the two spatial modes following the Knight’s jump pattern in chess, as shown in Fig. 2b. The deviations are smaller than when using the simpler strategy where only mode A is measured, as quantified by the simulation in Fig. 2 for different number of modes. Here, we chose for $p_{\text{test}}(r)$ the uniform distribution between $-0.5$ and $0.5$. While the difference is large for smaller interferometers, it reduces sharply for larger interferometers, and larger values of input squeezing are required to recover the advantage. The reasons for this behaviour are unclear and might be a subject for further research.
We suspect, however, that an increased expressibility comes from denser correlations between the modes that are measured and the modes that are used as output modes. These investigations are beyond the scope of this article.
III Experimental demonstration
@@
Having introduced the concept of implementing a large squeezed-light beam splitter network via measurement-induced interaction, we now experimentally demonstrate it using the setup in Fig. 1d. We employ a configuration with two cascaded asymmetric fiber interferometers, generating a 2D graph state within two optical spatial modes (denoted A and B) [4], with the first interferometer introducing a delay of $K_{1}=1$ clock cycle and the second interferometer a delay of $K_{2}=8$ clock cycles. We inject squeezed vacuum states generated by optical parametric oscillators (OPOs) into modes A and B with, respectively, 2.3 dB and 3 dB of squeezing in conjugate quadratures [22].
@@
The squeezed states are generated in the continuous wave regime, but the temporal structure introduced by the delay lines effectively defines time bins of duration 246.9 ns, corresponding to the $\sim$50 m length of the first fiber delay. Within these time bins, we define an anti-symmetric temporal mode that matches the OPO bandwidth but filters out noisy frequencies close to DC [4]. We detect both modes A and B with homodyne detectors: Mode A is measured in a variable basis $\vec{\theta}$, while the resulting measurement-induced graph state in mode B is characterized by $x$ (position) and $p$ (momentum) quadrature measurements. The 2D cluster state, ideally, extends infinitely. However, to generate finite parametrized states, we “chop” the 2D cluster using $x$-quadrature measurements, which remove correlations. Given the Gaussian nature of the state, full characterization is achieved by determining its multimode covariance matrix, which involves measuring all correlations for complete direct determination. To limit the number of measurements, we apply prior information about the state’s alignment in phase space and leverage the expected symmetry of the matrix, thereby reducing the number of required measurements considerably (for details, see [22]).
We first implement a programmable six-mode squeezed light interferometer by measurements on spatial mode A of a 12-mode graph state, which was in turn obtained through deletions by $x$-quadrature measurements of neighbouring modes of the larger 2D cluster state [23]. To illustrate the impact of different measurement basis choices, we carry out five different realizations in which a single homodyne angle, $\theta_{3}$, is varied in equal steps between $-\pi/2$ and $\pi/2$, as shown in Fig. 3a. For each of the realized output states, we characterize the covariance matrix with the results presented in Fig. 3b. Now, using Williamson and Bloch-Messiah-Euler decomposition of the covariance matrix, we construct the corresponding effective input squeezing and the unitary matrix of the induced interferometer, the amplitude and phase elements of which are illustrated in Fig. 3c and Fig. 3d for each of the five realizations. It is evident that altering just a single measurement angle has a significant impact on the configuration of the squeezed light interferometer, clearly demonstrating the ability to tune the quantum correlations of the resulting output.
@@
Figure 3: Measurement induced states and corresponding interferometers: five parametrized states and induced unitaries generated as a result of varying measurement parameters on temporal mode 3 of spatial mode B on a 12 mode (6 temporal modes in A and 6 temporal modes in B) 2D cluster state, demonstrating the ability of the method to induce parametrized interferometers. The respective five measurement settings (a) applied to 6 temporal modes in one of the spatial modes to obtain the induced graphs on the other spatial mode are used to generate the parametrized states and unitaries. Note that the third temporal mode undergoes varying measurement bases and the measurement setting for this mode is highlighted with a green overtone. The covariance matrices of the parametrized states (b) thus produced are decomposed to form the unitary matrix whose amplitude (c) and phase elements (d) are shown indexed by the mode numbers of the output modes.
As a next step, we consider a 400-mode squeezed light interferometer, representing a realization that is practically difficult to achieve using a conventional bulk or photonic integrated interferometer. Similar to the six-mode realization, we prepare the output state through the measurements of mode A, but now using 400 different measurement angles, $\vec{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{400})$, and reconstruct the covariance matrix through 24 measurement patterns across $x$ and $p$ quadratures [22]. To implement a large random unitary matrix, the 400 angles are randomly selected from the interval $[-\pi/2,\pi/2]$, and the resulting covariance matrix of size $800\times 800$ is characterized by the homodyne detector in mode B. Using this result, along with its decompositions, we deduce the input distribution of the squeezed states, as well as the unitary matrix representing the beam splitter transformation. We first show the amplitudes and the phases of the random unitary generated in the experiment in Fig. 4a. and Fig. 4b. Due to the chopping of the 2D cluster to a finite size ($800$ modes before the random measurement of $400$ modes), we see some edge effects in the matrix of amplitudes.
In Fig. 4, we also plot the amplitude and phase distributions of the elements of the resulting unitary matrix and compare them to the theoretical distribution for random unitaries as well as the distribution of numerically generated random matrices, which follow the distribution $P(a)=2(M-1)(1-a^{2})^{M-2}a$ [9, 10, 24] for the amplitude $a$ and a uniform distribution for the phase $\phi$. Like previous efforts [9, 10] we use fidelity to quantify how closely the distribution of the experimentally produced matrix elements matches that of unitaries drawn according to the Haar measure. The fidelity is defined by $F=\sum_{i}\sqrt{p_{i}q_{i}}$, where $p_{i}$ and $q_{i}$ represents the frequencies of the element amplitudes and phases of the experimentally produced matrices and the matrices sampled from the Haar measure. The calculated fidelities are $0.997$ and $0.999$ for the amplitude and phase elements, respectively. Such matrices often find applications in demonstrations of quantum advantage using photons[25, 26, 27, 28].
@@
Figure 4: Elements of the random matrix : a. Amplitude and, b. phase parts of the unitary that was induced through these random measurements. We note, in particular, the edge effects that appear in the matrix of amplitudes due to chopping of the 2D cluster to a finite size. c. The distribution of amplitude and, d. phase elements obtained in the experiment with respect to a instance of a numerically generated Haar random matrix and the theoretical prediction are shown here. We report fidelities of 0.997 and 0.999 in the cases of amplitudes and phases obtained, respectively, with respect to samples drawn from the Haar measure which is an improvement over the previous efforts.
IV Conclusions
@@
In this work, we have introduced and experimentally demonstrated a scalable, low-loss approach to generating multimode squeezed light interferometers based on measurement-induced operations. By leveraging quadrature measurements on time-domain graph states, we developed a flexible and programmable beam splitter network capable of accessing a wide class of unitaries. This method enables efficient parameter tuning through measurement, transforming a fixed, shallow circuit into a parameterized quantum circuit for generating parametric quantum states and interferometers. Notably, we demonstrated the realization of high-dimensional unitaries, including a random 400-mode interferometer with a distribution of elements that very closely resembles one whose samples are drawn from the Haar measure.
@@
Our approach aligns with the paradigm of parameterized quantum circuits used in discrete-variable systems, where fixed circuit structures are parameterized and iteratively tuned to optimize specific tasks. Similarly, our method enables precise tuning of a fixed graph state circuit through measurement-based parameters. Although methods and metrics for generating parametric states have been extensively studied in discrete-variable systems [21, 29, 30, 31], such an approach has been largely unexplored in continuous-variable systems.
@@
The ability to efficiently generate and tune high-dimensional quantum states and interferometers has far-reaching potential in quantum computational tasks, including achieving quantum advantage [25, 26, 27], simulating complex physical systems, and potentially addressing real-world challenges such as molecular modeling and drug discovery [32, 33, 34, 35, 36]. This development demonstrates a scalable platform for continuous-variable quantum technologies, positioning parameterized measurement-based circuits as a potentially useful tool for advancing the capabilities of NISQ systems.
@@
Although the current implementation is limited in producing exact states and interferometers for very large mode numbers, it demonstrates a pathway for further developments. By incorporating more advanced resource states and different initial graph states, refined measurement strategies and feedback schemes, as well as leveraging optimization techniques in order to produce reasonably close states to states of any particular interest, the range of achievable states and unitaries can be expanded, enhancing the practical applicability of the approach.

### v3

@@ III Expressibility
The expressibility of the parametrized circuits depends on several factors, which we explore in the following through simulation: 1) the architecture of the initial graph (i.e. the correlation complexity), 2) the number of modes, 3) the measurement strategy, and 4) the degree of input squeezing. Regarding 1), the complexity of the graph state correlations can be varied by increasing the number of cascaded interferometers in the setup as shown in Fig. 1. However, for this simulation we fix the initial graph as a 1D cluster state, that is, with only one delay line, $K_{1}=1$. As for 2), the number of modes in the state can easily be changed in the time domain. Here, we compare $N=8,16,32$. Regarding 3), the measurement strategy, we have so far assumed that spatial mode A was measured (as illustrated in Fig. 2a), thereby transforming the nodes in mode B. However, alternative measurement schemes can be employed, considering that with a fast optical switch each spatio-temporal mode could be directed either to the variable quadrature measurement or to the output. One should expect that increasing the number of circuit parameters, i.e. the length of $\vec{\theta}$ relative to $N$, should increase the expressibility. We explore that by comparing the previous “linear” measurement strategy with a new “Knight’s distance” (KD) measurement strategy, where the output modes alternate between the two spatial modes following the Knight’s jump pattern in chess, as shown in Fig. 2b. Finally, as for 4), the input squeezing $\vec{r}_{\text{in}}$ strongly affects both the accessible induced circuits $\mathcal{P}$ and the reference ensemble $\mathcal{H}$, given the method for constructing this outlined above. Furthermore, the induced (output) squeezing levels $\vec{r}_{\text{eff}}$ also depend on the measurement strategy: Additional measurements, as in the KD strategy, reduce the resulting output energy. To compare the two measurement strategies on even terms, we therefore choose the input squeezing levels such that the average photon number per mode, $\bar{n}_{\text{pm}}$, is identical. Further, all input modes are squeezed the same. For each $N$, we compare three output energy levels, $\bar{n}_{\text{pm}}=0.1,0.3,0.5$, which is obtained by choosing different squeezing levels for the different measurement strategies. For instance, for $N=8$, the squeezing levels are chosen to be $r_{\text{in}}=0.457,0.736,0.900$ for the linear pattern and $r_{\text{in}}=0.764,1.036,1.196$ for the KD pattern to obtain these $\bar{n}_{\text{pm}}$ values.
We present the result of these simulations in Fig. 2. For each configuration, we sampled 1000 circuits by sampling each homodyne phase angle $\theta_{i}$ uniformly between $\pi/2$ and $\pi/2$. For the reference ensemble, we similarly sampled 1000 unitaries $W$ and combined each of them with one of the induced squeezings $S_{\text{sq}}$ / $\vec{r}_{\text{eff}}$. All $10^{6}$ pairwise fidelities for each of the three terms in eq. (3) were then calculated, and the distributions for the two intra-set terms are plotted as histograms in panels (c-e): solid curves for the $\mathcal{P}$ distributions obtained with the linear pattern and dashed for the KD pattern, while the $\mathcal{H}$ distribution is filled. The deviations (3) representing expressibility are presented in panels (f-h). We see that, as the mean photon number per mode increases, the fidelities accumulate closer to 0. This effect is especially pronounced as $N$ increases. Looking at the deviation metric, there is not a definite trend, but in general, the deviations get lower as the squeezing/photon number increases. Furthermore, as $N$ increases, the weights of tails of the distributions due to the different measurement strategies also get closer to that of the reference distributions. This is reflected in the higher order moments of the different measurement strategies. Lastly, the KD states show a significantly higher overlap with the references as compared to the linear states across all values of $N$ and mean per-mode photon numbers. This is also clearly reflected in the values of the deviations across the different values of $t$, which are considerably lower for the KD strategy states than for states from the linear strategies, showing that a more advanced measurement strategy can indeed be highly beneficial. A more detailed argument for the reasons behind this is given in Supplementary Material, sec. 1 [supp].
IV Experimental demonstration
@@
We now experimentally demonstrate the measurement-induced squeezed-light network using the setup in Fig. 1d. We employ a configuration with two cascaded asymmetric fiber interferometers, generating a 2D graph state within two optical spatial modes (denoted A and B), with the first interferometer introducing a delay of $K_{1}=1$ clock cycle and the second interferometer a delay of $K_{2}=8$ clock cycles. We inject squeezed vacuum states generated by optical parametric oscillators (OPOs) into modes A and B with, respectively, 2.3 dB and 3.0 dB of squeezing in conjugate quadratures [supp], i.e. keeping $\vec{r}_{\text{in}}$ fixed.
@@
The squeezed states are generated in the continuous wave regime, but the temporal structure introduced by the delay lines effectively defines time bins of duration 246.9 ns, corresponding to the $\sim$50 m length of the first fiber delay. Within these time bins, we define an anti-symmetric temporal mode that matches the OPO bandwidth but filters out noisy frequencies close to DC [cluster19]. We detect both modes A and B with homodyne detectors: Mode A is measured in a variable basis $\vec{\theta}$, while the resulting measurement-induced graph state in mode B is characterized by $x$ (position) and $p$ (momentum) quadrature measurements. The 2D cluster state, ideally, extends infinitely. However, to generate finite parametrized states, we “chop” the 2D cluster using $x$-quadrature measurements, which remove correlations. Given the Gaussian nature of the state, full characterization is achieved by determining its multimode covariance matrix, which involves measuring all correlations for complete direct determination. To limit the number of measurements, we leverage the expected symmetry of the covariance matrix, thereby reducing the number of required measurements considerably (for details, see [supp]).
@@
Figure 3: Measurement induced states and corresponding interferometers: five parametrized states and their induced unitaries generated as a result of varying measurement parameters on temporal mode 3 of spatial mode A on a 12 mode (6 temporal modes in A and 6 temporal modes in B) cluster state, demonstrating the ability of the method to induce parametrized interferometers. (a) The respective five different measurement settings applied to the 6 temporal modes. Note that the third temporal mode undergoes varying measurement bases, as highlighted in green. The covariance matrices of the parametrized states thus produced (b) are decomposed to form the unitary matrix whose amplitude (d) and phase elements (e) are shown indexed by the mode numbers of the output modes while the effective induced squeezing that enters the induced interferometers in also shown (c).
We first implement a programmable six-mode squeezed light interferometer by measurements on spatial mode A of a 12-mode graph state, which was in turn obtained through deletions by $x$-quadrature measurements of neighbouring modes of the larger 2D cluster state [GraphicalCalc]. To illustrate the impact of different measurement basis choices, we carry out five different realizations in which a single homodyne angle, $\theta_{3}$, is varied in equal steps between $\pi/2$ and $\pi/2$, as shown in Fig. 3a. For each of the realized output states, we present in Fig. 3b the covariance matrix with the thermalization from random measurement outcomes removed as described above and in [supp], sec. 3.6. The states are generated with relatively high fidelity of 0.91, 0.93, 0.89, 0.91 and 0.88, respectively, with respect to the simulated states with the input squeezing equal to the mean of squeezing values extracted from the experimental data (Supplementary Information, section 3.1.2). Most of the loss of fidelity can be attributed to loss and unequal squeezing of the two OPOs used for the generation of the cluster state. Now, using Williamson and Bloch-Messiah-Euler decomposition of the covariance matrix, we construct the corresponding effective input squeezing and the unitary matrix of the induced interferometer, the amplitude and phase elements of which are illustrated in Fig. 3c and Fig. 3d for each of the five realizations. It is evident that altering just a single measurement angle has a significant impact on the configuration of the squeezed light interferometer, clearly demonstrating the ability to tune the quantum correlations of the resulting output.
@@
Figure 4: Elements of the random matrix : (a) Amplitude and, (b) phase parts of the unitary that was induced through these random measurements. (c) The probability distribution of amplitude and, (d) phase elements obtained in the experiment with respect to a instance of a numerically generated Haar random matrix and the theoretical prediction are shown here. We report fidelities of 0.997 and 0.999 in the cases of amplitudes and phases obtained, respectively, with respect to samples drawn from the Haar measure which is an improvement over the previous efforts.
As a next step, we consider a 400-mode squeezed light interferometer, representing a realization that is practically difficult to achieve using a conventional bulk or photonic integrated interferometer. Similar to the six-mode realization, we prepare the output state through the measurements of mode A, but now using 400 different measurement angles, $\vec{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{400})$, and reconstruct the covariance matrix through 22 measurement patterns across $x$ and $p$ quadratures ([supp], sec. 3.4). To implement a large random unitary matrix, the 400 angles are sampled uniformly from the interval $\pi/2,\pi/2]$, and the resulting covariance matrix of size $800\times 800$ is characterized by the homodyne detector in mode B. Using this result, along with its decompositions, we deduce the distribution of the induced squeezed states, as well as the unitary matrix representing the beam splitter transformation. We first show the amplitudes and the phases of the random unitary generated in the experiment in Fig. 4a and Fig. 4b.
In Fig. 4, we also plot the amplitude and phase distributions of the elements of the resulting unitary matrix and compare them to the theoretical distribution for Haar-random unitaries as well as the distribution of numerically generated random matrices, which follow the distribution $P(a)=2(N-1)(1-a^{2})^{N-2}a$ [USTC20, Xanadu21, LargeHaar] for the amplitude $a$ and a uniform distribution for the phase $\phi$.
From Fig. 4a, the unitary matrix exhibits a bit of structure, making it not completely random. To test the randomness, we show the distribution of eigenvalues and eigenphases of the matrices generated in supplementary information, section 3.7 (Figure 13, 14). We find the results consistent with that of a circular unitary ensemble. Furthermore, like previous efforts [USTC20, Xanadu21] we use fidelity to quantify how closely the distribution of the experimentally produced matrix elements matches that of unitaries drawn according to the Haar measure. The fidelity is defined by $F=\sum_{i}\sqrt{p_{i}q_{i}}$, where $p_{i}$ and $q_{i}$ represents the probability densities of the element amplitudes and phases of the experimentally produced matrices and the matrices sampled from the Haar measure. The calculated fidelities are $0.997$ and $0.999$ for the amplitude and phase elements, respectively. Such matrices often find applications in demonstrations of quantum advantage using photons[QAdv1, QAdv2, QAdv3, Cleo].
V Conclusions
@@
In this work, we have introduced and experimentally demonstrated a scalable, low-loss approach to generating multimode squeezed light interferometers based on measurement-induced operations. By leveraging quadrature measurements on time-domain graph states, we developed a flexible and programmable beam splitter network capable of accessing a wide class of unitaries. This method enables efficient parameter tuning through measurement, transforming a fixed, shallow circuit into a parametrized quantum circuit for generating parametric quantum states and interferometers. Notably, we demonstrated the realization of high-dimensional unitaries, including a random 400-mode interferometer with a distribution of elements that very closely resembles one whose samples are drawn from the Haar measure.
@@
Our approach aligns with the paradigm of parametrized quantum circuits used in discrete-variable systems, where fixed circuit structures are parametrized and iteratively tuned to optimize specific tasks. Similarly, our method enables precise tuning of a fixed graph state circuit through measurement-based parameters. Although methods and metrics for generating parametric states have been extensively studied in discrete-variable systems [PQC1, PQC2, PQC3, PQC4], such an approach has been largely unexplored in continuous-variable systems.
@@
The ability to efficiently generate and tune high-dimensional quantum states and interferometers has far-reaching potential in quantum computational tasks, including achieving quantum advantage [QAdv1, QAdv2, QAdv3], simulating complex physical systems, and potentially addressing real-world challenges such as molecular modeling and drug discovery [Amb1, Amb2, Amb3, DD1, DD2]. This development demonstrates a scalable platform for continuous-variable quantum technologies, positioning parametrized measurement-based circuits as a potentially useful tool for advancing the capabilities of NISQ systems.
@@
Although the current implementation is limited in producing exact states and interferometers for very large mode numbers, it demonstrates a pathway for further developments. By incorporating more advanced resource states and different initial graph states, refined measurement strategies and feedback schemes, as well as leveraging optimization techniques in order to produce states reasonably close to states of any particular interest, the range of achievable states and unitaries can be expanded, enhancing the practical applicability of the approach.
VI Acknowledgements
We acknowledge support from the Danish National Research Foundation (bigQ, DNRF0142), EU ERC project ClusterQ (grant agreement no. 101055224, ERC-2021-ADG), Innovation Fund Denmark (PhotoQ, 3155-00024A, and QuantERA - ClusSTAR, 3155-00024A), and EU (CLUSTEC, grant agreement no. 101080173, and QuantERA - ClusSTAR).